OpenStudy (anonymous):

(y+6)^2/y^2-36 X 3y-18/2y +12

OpenStudy (nubeer):

ok first u have to make factors of that y^2-36 in the denominator..

OpenStudy (anonymous):

(y+6)(y-6)

OpenStudy (nubeer):

ok good now u can cancel y+6 with the one in numerator..

OpenStudy (anonymous):

so now it's (y+6)/(y-6)

OpenStudy (nubeer):

yes.. on other side.. u can take 3 common from numerator and 2 from denominator.

OpenStudy (anonymous):

I don't get how to do that part

OpenStudy (nubeer):

take 3 common in numerator. 3(y-6) and now denominator. 2(y+6)

OpenStudy (anonymous):

now do you cancel more?

OpenStudy (nubeer):

yes if possible

OpenStudy (anonymous):

wouldn't it just be 3/2?

OpenStudy (nubeer):

\[\frac{ y+6 }{ y-6 } * \frac{ 3( y-6)}{ 2(y+6) }\]

OpenStudy (nubeer):

yes.

OpenStudy (anonymous):

is that right? i feel like you can't cancel them if the coefficients are there in the second pa\rt

OpenStudy (austinl):

I realize that this has almost been solved but. \[\frac{(y+6)^2}{y^2-36}\times \frac{(3y-18)}{(2y+12)}\] \[\frac{(3y-18)(y+6)}{(y-6)(2y+12)}\] \[\frac{3y^2+18y-18y}{2y^2+12y-12y}\] \[\frac{3y^2}{2y^2}\] Then, you can cancel, and you get your answer.

OpenStudy (austinl):

I of course skipped a step at the beginning, I didn't show my factoring.

OpenStudy (anonymous):

so is it 3/2 or do you keep the y^2

OpenStudy (austinl):

There is a y^2 on top and bottom, so you could cancel them. The answer should be 3/2.

OpenStudy (anonymous):

could you help me with a division one i'll ask

OpenStudy (austinl):

Sure!